OpenStudy (laweejay):

Find the point on the parabola y = (x-3)^2 that is closest to the origin...

OpenStudy (dumbcow):

distance from origin \[d = \sqrt{x^2 +y^2}\] minimize distance by setting derivative equal to 0 \[\frac{2x +2y \frac{dy}{dx}}{2 \sqrt{x^2+y^2}} = 0\] \[\rightarrow \frac{dy}{dx} = -\frac{x}{y}\] \[2(x-3) = -\frac{x}{(x-3)^2}\] \[2x^3 -18x^2 +55x -54=0\] solve for x